OpenStudy (anonymous):

use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions. 1. s^2 - 4s + 4 = 0 2. 5/6x^2 - 7x - 6/5 = 0 3. 7a^2 + 8a + 2 = 0 4. (7)^1/2y^2 - 6y - 13(7)^1/2 = 0

OpenStudy (saifoo.khan):

Mini for rescue!

OpenStudy (mimi_x3):

The discriminant is : D= b^2-4ac if D greater than or equal to 0, the roots are real If D is less than 0 then the roots are not real If D=0 the roots are equal If D is a perfect square, the roots are rational

OpenStudy (anonymous):

ii failed this 4 times.

OpenStudy (mimi_x3):

SUch a failure "supaman"

OpenStudy (anonymous):

lmao I'm a superman when it comes to woman. lol NOT math

OpenStudy (mimi_x3):

lols, your C like Saifi :P You should change your username then to failman, souds better :P

OpenStudy (anonymous):

LOL. naw mami. I shall not, but you should help me [:

OpenStudy (mimi_x3):

mami? I just helped you, you should do it yourself, to stop being a failure

OpenStudy (anonymous):

LOL i failed it 4 times, I wouldn't be on here if i could do it myself. Don't I have to like do the work or just point out if its real or not?

OpenStudy (mimi_x3):

I just told you, use he discriminant then decide. s^2 -4s + 4 = 0 a = 1 , b=-4 , c = 4 then substitute it into the discriminant b^2-4ac

OpenStudy (anonymous):

wtf, i never done discriminant.

OpenStudy (anonymous):

ill just pay for a tutor. thanks much.

OpenStudy (saifoo.khan):

i can teach! ^_^

OpenStudy (saifoo.khan):

want to study? free. :p

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

if you want too, if you can make since that is.

OpenStudy (mimi_x3):

lols, for free?! then you should help me then :P

OpenStudy (mimi_x3):

Im more worthwhile to help :P

OpenStudy (anonymous):

LOL mimi i can help you with other needs [:

OpenStudy (anonymous):

like english? LOL

OpenStudy (saifoo.khan):

The discriminant is : D= b^2-4ac if D greater than or equal to 0, the roots are real If D is less than 0 then the roots are not real If D=0 the roots are equal so for the first one: s^2 - 4s + 4 = 0 b^2 - 4ac \[(-4)^2 - 4(1)(4)\]\[16-16\]ZERO. as the result we are getting is zero, so this tells us that the graph will cute the x-axis once. and both the roots of the x will be same.. or in other words perfect square.

OpenStudy (saifoo.khan):

getting it?

OpenStudy (mimi_x3):

Plagiarism DETECTED !

OpenStudy (anonymous):

Kind of, The main part I'm stuck on is where you 4ac and so forth.

OpenStudy (saifoo.khan):

what grade r u in?

OpenStudy (saifoo.khan):

http://www.twiddla.com/657721

OpenStudy (saifoo.khan):

come here.

OpenStudy (mimi_x3):

"will cute the axis" lmao

OpenStudy (anonymous):

I'm in college math.

OpenStudy (mimi_x3):

College!? :O Thats year 10 hehe :P

OpenStudy (saifoo.khan):

Plagrism confirmed @mini.. @superman, i teach college algebra.

OpenStudy (mimi_x3):

here*

OpenStudy (anonymous):

lol mmi i seen that too. but its okay I understood.

OpenStudy (mimi_x3):

Saifi, you NEVER help me, argh :P

OpenStudy (saifoo.khan):

Are u in UoP?

OpenStudy (saifoo.khan):

Mini what type of help do u want my friend?

OpenStudy (anonymous):

dude i dont know the difference, I HATE MATH. unless its pre algebra and algebra in high school just for some reason i suck in this.

OpenStudy (saifoo.khan):

are u on that link?

OpenStudy (mimi_x3):

Geometrical Properties of Parabola, it's killing me

OpenStudy (mimi_x3):

Algebra's easy, geometry sucks

OpenStudy (saifoo.khan):

Both are easy.

OpenStudy (anonymous):

http://www.twiddla.com/657721 go to the link mimi